Are young infants more likely to imitate actions that are modeled by a person or simulated by an object? This question was the basis of a research study summarized in the article “The Role of Person and Object in Eliciting Early Imitation." One action examined was mouth opening. This action was modeled repeatedly by either a person or a doll, and the number of times that the infant imitated the behavior was recorded. 80 infants participated, with 35 exposed to a human model and 45 exposed to the doll. The personal model had sample mean of 5.14 and sample standard deviation of 1.60, while the doll model had a sample mean of 3.46 and sample standard deviation of 1.30. Is there sufficient evidence to conclude with a 5% significance level that the mean number of imitations is different between infants who watch a human model and infants who watch a doll? Assume equal population variance between infant groups.
Answer #1

Ans:

Pooled standard deviation=sqrt(((35-1)*1.6^2+(45-1)*1.3^2)/(35+45-2))=1.4385

standard error=1.4385*sqrt((1/35)+(1/45))=0.3242

Test statistic:

t=(5.14-3.46)/0.3242

t=5.182

df=35+45-2=78

p-value=tdist(5.182,78,2)=0.0000

As,p-value<0.05,we reject the null hypothesis.

There is sufficient evidence to conclude with a 5% significance level that the mean number of imitations is different between infants who watch a human model and infants who watch a doll